# On the connection between Bochner’s theorem on positive definite maps and Choi theorem on complete positivity

### v1

@@ III.2 Fourier Inversion for finite inverse semigroup
Let us consider the matrix unit inverse semigroup $S_{M}:=\{e_{ij}\}_{i,j=1}^{m}\cup\{0\}$. Then the Fourier inversion formula in Eq. (37) becomes the Choi inversion formula in Eq. (14).
Proof: Let us start by recalling that two elements $s$ and $t$ belonging to an inverse semigroup $S$ satisfies $s\leq t$ if and only if there exists an idempotent $e\in S$ such that $s=et$ (see Definition (D)). Trivially we have $0\leq e_{ij}$ for all $i,j\in\{1,2,...,m\}$. Now, let $e_{ij}\leq e_{kl}$. Then there exists an idempotent $e_{pp}$ such that $e_{ij}=e_{pp}e_{kl}=e_{pl}\delta_{pk}$. As $e_{ij}\neq 0$, we must have $p=k$ which implies that $e_{ij}=e_{kl}$, which further implies that the only element $x\in\{e_{ij}\}_{i,j=1}^{m}$ that satisfies $e_{ij}\leq x$ is $x=e_{ij}$. From Eq. (8), we have that the groupoid basis corresponding to $e_{ij}$ is $e_{ij}$ itself, i.e., $\lfloor e_{ij}\rfloor=e_{ij}$ and from Eq. (40), we have $\widetilde{\Phi}(\lfloor e_{ij}\rfloor)=\Phi(e_{ij})$. As $D_{1}$ contains $m$ idempotents, we have $r_{1}=m$ and $G_{e_{ii}}$ being a group with only one element in it, we have $|G_{e_{ii}}|=1$. Using all these information in Eq. (37), we have
@@
Now, up to equivalence, there is exactly one irreducible representation of the full matrix algebra $M_{m}(\mathbb{C})$ [16]. We can “take” this representation to be the identity map $\mathrm{id}$" on the matrix algebra $M_{m}(\mathbb{C})$ (under the identification of $M_{m}(\mathbb{C})$ with $\mathrm{End}(\mathbb{C}^{m})$). Now using the fact that the dimension of this representation is $m$, i.e., $d_{\mathrm{id}}=m$ and the Fourier transform of the map $\Phi$ with respect to the representation $\mathrm{id}$" is the Choi matrix of $\Phi$, i.e., $\widehat{\Phi}(\mathrm{id})=C_{\Phi}$, we have
$\Phi(e_{ij})=\operatorname{tr}_{m}\big{[}(e_{ji}\otimes\mathbb{I})C_{\Phi}\big{]},$ (57)
or equivalently for an arbitrary $x\in M_{n}(\mathbb{C})$, we have
@@
where $\tau$ represents the transposition map. This completes the proof. $\Box$
@@
The Plancherel Formula is given by
@@
where $r_{k}$ and $G_{e_{k}}$ are defined in Theorem (2).
Proof: Form Eq. (37), we have
@@
Notice that $r_{k}$ and $|G_{e_{k}}|$ does not change as $s\in D_{k}$ if and only if $s^{-1}\in D_{k}$. Now multiplying both the sides by $\widetilde{\Psi}(\lfloor s\rfloor)$ and summing over $s$, we have
@@
For matrix unit semigroup $S_{M}$, we have $r_{k}=m$, $|G_{e_{k}}|=1$, $\widetilde{\Phi}(\lfloor e_{ij}\rfloor)=\Phi(e_{ij})$ and $\widetilde{\Psi}(\lfloor e_{ij}\rfloor)=\Psi(e_{ij})$. Hence, the above formula reduces to
@@
III.3 Positive definite maps and Bochner’s theorem for finite inverse semigroup
Let $G$ be a group, and $\mathcal{H}=\mathbb{C}^{n}$ be a Hilbert space. A map $\Phi:G\to\mathcal{B}(\mathcal{H})$ is called positive definite if the matrix $\Phi(g^{-1}g^{\prime})]\in\mathcal{B}(\mathcal{H}^{\bigoplus|G|})$ is positive semidefinite, i.e.,
@@ III.3 Positive definite maps and Bochner’s theorem for finite inverse semigroup
for all subset $h:=\{h_{s}\}_{s\in G}\subset\mathbb{C}^{n}$ whose elements are indexed by the elements of $G$. The positive definite-ness of a maps on a group is related to the positive semidefinite-ness of its Fourier transform by Bochner’s theorem. Let $G$ be a finite group and $f:G\to\mathbb{C}$ be a complex valued function. The Bochner’s theorem states that $f$ is positive definite if and only if its Fourier transform $\widehat{f}(\rho)$ is positive semidefinite for all $\rho\in\mathrm{IrrU(G)}$, where $\mathrm{IrrU(G)}$ is the complete set of inequivalent irreducible unitary representations of the group $G$. The proof can be found in[17]. For matrix valued map $\Phi:G\to M_{n}(\mathbb{C})$, Bochner’s theorem becomes the following.
@@
We provide a proof in Appendix (D).
Now we are in a position to formulate and prove Bochner’s theorem for finite inverse semigroup. We begin by defining positive definite maps on the contracted algebra of a finite inverse semigroup.
@@
Let $S$ be a finite inverse semigroup with $z$ being the zero element. A linear map $\Phi:\mathbb{C}_{0}[S]\to\mathcal{B}(\mathcal{H})$ where $\mathcal{H}=\mathbb{C}^{n}$, is called positive definite if the matrix $\Phi(s^{-1}s^{\prime})]\in\mathcal{B}(\mathcal{H}^{\bigoplus(|S|-1)})$ is positive semidefinite, i.e.,
@@
for all subset $h:=\{h_{s}\}_{s\in S\setminus\{0\}}\subset\mathbb{C}^{n}$ whose elements are indexed by the nonzero elements of $S$.
The following propositions will be useful to prove our main results.
@@
for all subset $h:=\{h_{s}\}_{s\in S\setminus\{0\}}\subset\mathbb{C}^{n}$ whose elements are indexed by the nonzero elements of $S$.
Proof: We start by recalling the definition of the groupoid basis $\lfloor s\rfloor$ of $\mathbb{C}_{0}[S]$ (see Eq. (8)).
@@
where $t\neq 0$. The above equation implies that
$\lfloor s^{-1}\rfloor$ $=\sum_{\begin{subarray}{c}t\in S\\ t\leq s\end{subarray}}\mu(t^{-1},s^{-1})t^{-1}=\sum_{\begin{subarray}{c}t\in S\\ t\leq s\end{subarray}}\mu(t^{-1}t,s^{-1}s)t^{-1}=\sum_{\begin{subarray}{c}t\in S\\ t\leq s\end{subarray}}\mu(t,s)t^{-1},$ (65)
where in the last two equalities we have used the fact that $\mu(t,s)=\mu(tt^{-1},ss^{-1})=\mu(t^{-1}t,s^{-1}s)$ (See [18]). Now, with $s,s^{\prime}\neq 0$, we have
@@
Defining the set $h^{0}:=\{h^{0}_{t}\}$ consisting of elements $h^{0}_{t}:=\sum_{\begin{subarray}{c}s\in S\\ t\leq s\end{subarray}}\mu(t,s)h(s)$ (which can be inverted as $h_{s}=\sum_{\begin{subarray}{c}t\in S\\ t\geq s\end{subarray}}h^{0}_{t}$), we can rewrite the above equation as
@@
for all $D_{k}$ and all subset $h:=\{h_{s}\}_{s\in D_{k}}\subset\mathbb{C}^{n}$ whose elements are indexed by the elements of $D_{k}$.
Proof: The multiplication rule for the groupoid basis as given by Eq. (10) makes the matrix $\Phi(\lfloor s^{-1}\rfloor\lfloor s^{\prime}\rfloor)]$ block diagonal because $\lfloor s^{-1}\rfloor\lfloor s^{\prime}\rfloor=0$ when $ss^{-1}\neq s^{\prime}s^{\prime-1}$. Now, if $ss^{-1}=s^{\prime}s^{\prime-1}$ then $s$ and $s^{\prime}$ belong to the same $\mathcal{D}$-class which implies that if $s$ and $s^{\prime}$ belong to different $\mathcal{D}$-classes then $ss^{-1}\neq s^{\prime}s^{\prime-1}$. Hence $\Phi$ is positive definite if and only if for all $\mathcal{D}$-classes $D_{k}$ the following relation holds:
@@
for all subset $h:=\{h_{s}\}_{s\in D_{k}}\subset\mathbb{C}^{n}$ whose elements are indexed by the elements of $D_{k}$. This completes the proof. $\Box$
Proposition 5.
(Schur Orthogonality relation for inverse semigroup:) Let $S$ be a finite inverse semigroup and $\{D_{k}\}_{k=1}^{n}$ represents its $\mathcal{D}$-classes. Let $s\in D_{k}$ and $\overline{\rho}_{k}\in\mathrm{IrrU}(\mathbb{C}_{0}[S])$, where $\mathrm{IrrU}(\mathbb{C}_{0}[S])$ represents a complete set of inequivalent irreducible representations of $\mathbb{C}_{0}[S]$ induced by the complete set of inequivalent irreducible unitary representations of the maximal subgroups of $S$ as described in Theorem (J) . Let $a,b,c,d\in D_{k}$ be idempotents. Then we have
@@
where $\overline{((\overline{\rho}_{k}(\lfloor s\rfloor))_{a,b})_{i,j}}$ represents the complex conjugate of the $(i,j)$-th matrix element of the $(a,b)$-th block of $\overline{\rho}_{k}(\lfloor s\rfloor)$.
Proof: The action of the induced representation $\overline{\rho}_{k}$ on a groupoid basis is given by
@@
Notice that $u:=p^{-1}_{\mathrm{ran}(s)}\,s\,p_{\mathrm{dom}(s)}\in G_{e_{k}}$. Now from Eq. (72), it is clear that $\overline{((\overline{\rho}_{k}(\lfloor s\rfloor))_{a,b})_{i,j}}=\overline{(\rho_{k}(u))_{i,j}}\delta_{a,\mathrm{ran}(s)}\delta_{b,\mathrm{dom}(s)}$. Now,
$\sum_{s\in D_{k}}((\overline{\rho}_{k}(\lfloor s\rfloor))_{a,b})_{i,j}\overline{((\overline{\rho^{\prime}}_{k}(\lfloor s\rfloor))_{c,d})_{k,l}}$ $=\sum_{s\in D_{k}}(\rho_{k}(u))_{i,j}\delta_{a,\mathrm{ran}(s)}\delta_{b,\mathrm{dom}(s)}\overline{(\rho^{\prime}_{k}(u))_{k,l}}\delta_{c,\mathrm{ran}(s)}\delta_{d,\mathrm{dom}(s)}.$ (73)
Since the map $\vartheta:D_{k}\to M_{r_{K}}(G_{e_{k}})$ given by $s\mapsto(p^{-1}_{\mathrm{ran}(s)}\,s\,p_{\mathrm{dom}(s)})E_{\mathrm{ran}(s),\mathrm{dom}(s)}$ is one-to-one correspondence with the inverse given by $gE_{e,f}\mapsto p_{e}gp_{f}^{-1}$, any $s\in D_{k}$ can be written as $s=p_{e}gp_{f}^{-1}$ for some pair of idempotents $(e,f)\in D_{k}$ and $g\in G_{e_{k}}$. It should be noted that $\mathrm{ran}(p_{e}gp_{f}^{-1})=e$ and $\mathrm{dom}(p_{e}gp_{f}^{-1})=f$ for all $g\in G_{e_{k}}$. A necessary condition for the terms in the summation of Eq. (73) to be non-vanishing is $a=\mathrm{ran}(s)$ and $b=\mathrm{dom}(s)$. So the possible non-vanishing contributions come from summing over those $s\in D_{k}$ which has the form $s=p_{a}gp_{b}^{-1}$. With this observation, we have
@@
Now by applying the Schur orthogonality relation for the inequivalent irreducible unitary representations $\rho_{k}$ and $\rho^{\prime}_{k}$ in Eq. (74), and using the fact that $d_{\bar{\rho}_{k}}=r_{k}d_{\rho_{k}}$, where $d_{\bar{\rho}_{k}}$ represents the dimension of the induced representation $\overline{\rho}_{k}$, we have
@@
Before proceeding further, we recall from Eq. (38) that $\overline{\rho}_{k}(\lfloor s\rfloor)=E_{\mathrm{ran}(s),\mathrm{dom}(s)}\otimes\rho_{k}(p^{-1}_{\mathrm{ran}(s)}\,s\,p_{\mathrm{dom}(s)})$ and $\overline{\rho}_{k}(\lfloor s\rfloor)=0$ if $s\notin D_{k}$. So, we have $(\overline{\rho}_{k}(\lfloor s\rfloor))^{\dagger}=E_{\mathrm{dom}(s),\mathrm{ran}(s)}\otimes\rho_{k}(p^{-1}_{\mathrm{dom}(s)}\,s^{-1}\,p_{\mathrm{ran}(s)})$ and using Eq. (50), we have $(\overline{\rho}_{k}(\lfloor s\rfloor))^{\dagger}=\overline{\rho}_{k}(\lfloor s^{-1}\rfloor)$. This will be useful in proving Bochner’s theorem for finite inverse semigroup.
Theorem 4.
(Bochner’s theorem for finite inverse semigroups:) Let $S$ be a finite inverse semigroup and $\Phi:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ be a linear map. Then the linear map $\widetilde{\Phi}:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ defined by $\widetilde{\Phi}(\lfloor s\rfloor)=\sum_{\begin{subarray}{c}t\in S\\ t\geq s\end{subarray}}\Phi(t)$ is positive definite if and only if $\widehat{\Phi}(\overline{\rho}_{k})$ is positive semidefinite for all $\overline{\rho}_{k}\in\mathrm{IrrU}(\mathbb{C}_{0}[S])$, where $\mathrm{IrrU}(\mathbb{C}_{0}[S])$ represents a complete set of inequivalent irreducible representations of $\mathbb{C}_{0}[S]$ induced by the complete set of inequivalent irreducible unitary representations of the maximal subgroups of $S$.
Proof: We begin by recalling the inversion formula for finite inverse semigroup
$\widetilde{\Phi}(\lfloor s\rfloor)$ $=\frac{1}{r_{k}|G_{e_{k}}|}\sum_{\sigma\in\mathrm{IrrU}(\mathbb{C}_{0}[S])}d_{\sigma}\operatorname{tr}_{1}\big{[}(\sigma(\lfloor s^{-1}\rfloor)\otimes\mathbb{I})\widehat{\Phi}(\sigma)\big{]},$ (76)
where $s\in D_{k}$. Since $\sigma\in\mathrm{IrrU}(\mathbb{C}_{0}[S])$, we have $\sigma(\lfloor s^{-1}\rfloor)=(\sigma(\lfloor s\rfloor))^{\dagger}$. So the above equation can be written as
@@
Let $s,t\in D_{k}$ are such that $s^{-1}s=tt^{-1}$. For such a pairs of $s$ and $t$, Eq. (78) becomes

### v2

@@ 3.2 Fourier Inversion for Finite Inverse Semigroups
Let $S_{M}:=\{e_{ij}\}_{i,j=1}^{m}\cup\{0\}$ be the matrix unit inverse semigroup. Then the Fourier inversion formula in Eq. (46) reduces to the Choi inversion formula in Eq. (21).
Proof: Recall that for two elements $s,t$ in an inverse semigroup $S$, one has $s\leq t$ if and only if there exists an idempotent $e\in S$ such that $s=et$ (see Definition C). Clearly, $0\leq e_{ij}$ for all $i,j\in\{1,2,...,m\}$.
Now, suppose that $e_{ij}\leq e_{kl}$. Then there exists an idempotent $e_{pp}$ such that $e_{ij}=e_{pp}e_{kl}=e_{pl}\delta_{pk}$. Since $e_{ij}\neq 0$, it follows that $p=k$, and hence $e_{ij}=e_{kl}$. Therefore the only element $x\in\{e_{ij}\}_{i,j=1}^{m}$ satisfying $e_{ij}\leq x$ is $x=e_{ij}$ itself.
It follows from Eq. (8) that the groupoid basis corresponding to $e_{ij}$ is $e_{ij}$ itself, i.e., $\lfloor e_{ij}\rfloor=e_{ij}$. Consequently, by Eq. (49), we have $\widetilde{\Phi}(\lfloor e_{ij}\rfloor)=\Phi(e_{ij})$. Since the $\mathcal{D}$-class $D_{1}$ contains $m$ idempotents, we have $r_{1}=m$. Moreover, $G_{e_{ii}}$ being a group with only one element in it, we have $|G_{e_{ii}}|=1$.
Substituting these observations into Eq. (46), we obtain
@@
Now, up to equivalence, there is exactly one irreducible representation of the full matrix algebra $M_{m}(\mathbb{C})$ [39]. We can “take” this representation to be the identity map $\mathrm{id}$" on the matrix algebra $M_{m}(\mathbb{C})$ (under the identification of $M_{m}(\mathbb{C})$ with $\mathrm{End}(\mathbb{C}^{m})$). Now using the fact that the dimension of this representation is $m$, i.e., $d_{\mathrm{id}}=m$ and the Fourier transform of the map $\Phi$ with respect to the representation $\mathrm{id}$" is the Choi matrix of $\Phi$, i.e., $\widehat{\Phi}(\mathrm{id})=C_{\Phi}$, we obtain
$\Phi(e_{ij})=\operatorname{tr}_{m}\big[(e_{ji}\otimes\mathbb{I})C_{\Phi}\big].$ (66)
Equivalently, for an arbitrary $x\in M_{n}(\mathbb{C})$, we have
@@
where $\tau$ denotes the transpose map. This completes the proof. $\Box$
@@
The Plancherel formula for a finite inverse semigroup $S$ is given by
@@
where $r_{k}$ and $G_{e_{k}}$ are defined in Theorem 2.
Proof: Since $s\in D_{k}$ if and only if $s^{-1}\in D_{k}$, the quantities $r_{k}$ and $|G_{e_{k}}|$ remain unchanged. Consequently, form Eq. (46), we have
@@
Now multiplying both the sides by $\widetilde{\Psi}(\lfloor s\rfloor)$ and summing over $s$ with $s\neq 0$, we obtain
@@
For matrix unit inverse semigroup $S_{M}$, we have $r_{k}=m$, $|G_{e_{k}}|=1$, $\widetilde{\Phi}(\lfloor e_{ij}\rfloor)=\Phi(e_{ij})$ and $\widetilde{\Psi}(\lfloor e_{ij}\rfloor)=\Psi(e_{ij})$. In this case, the Plancherel formula reduces to
@@
3.3 Positive Definite Maps and Bochner’s theorem for Finite Inverse Semigroups
Let $G$ be a finite group, and let $\mathcal{H}=\mathbb{C}^{n}$ be a Hilbert space. A map $\Phi:G\to\mathcal{B}(\mathcal{H})$ is called positive definite if the matrix $\Phi(g^{-1}g^{\prime})]\in\mathcal{B}(\mathcal{H}^{\bigoplus|G|})$ is positive semidefinite [40], i.e.,
@@ 3.3 Positive Definite Maps and Bochner’s theorem for Finite Inverse Semigroups
for all subsets $h:=\{h_{s}\}_{s\in G}\subset\mathbb{C}^{n}$ whose elements are indexed by the elements of $G$. The positive definiteness of maps on a group is closely related to the positivity of their Fourier transforms via Bochner’s theorem. Let $G$ be a finite group, and let $f:G\to\mathbb{C}$ be a complex valued function. Bochner’s theorem states that $f$ is positive definite if and only if its Fourier transform $\widehat{f}(\rho)$ is positive semidefinite for all $\rho\in\mathrm{IrrU(G)}$, where $\mathrm{IrrU(G)}$ is the complete set of inequivalent irreducible unitary representations of the group $G$. A proof can be found in [41]. For matrix valued map $\Phi:G\to M_{n}(\mathbb{C})$, the above Bochner’s theorem remains valid and can be stated explicitly as follows:
@@
A proof is provided in Appendix C.
Now we are in a position to formulate and prove Bochner’s theorem for finite inverse semigroups. We begin by defining positive definite maps on the contracted algebra of a finite inverse semigroup.
@@
Let $S$ be a finite inverse semigroup with 0 denoting the zero element, and let $\mathcal{H}=\mathbb{C}^{n}$ be a Hilbert space. A linear map $\Phi:\mathbb{C}_{0}[S]\to\mathcal{B}(\mathcal{H})$ is called positive definite if the matrix $\Phi(s^{-1}s^{\prime})]\in\mathcal{B}(\mathcal{H}^{\bigoplus(|S|-1)})$ is positive semidefinite, i.e.,
@@
for all subsets $h:=\{h_{s}\}_{s\in S\setminus\{0\}}\subset\mathbb{C}^{n}$ whose elements are indexed by the nonzero elements of $S$.
In the following, we present two alternative characterizations of positive definite maps (Proposition 3 and Proposition 4), which will be used in the proof of Bochner’s theorem for finite inverse semigroups.
@@
for all subsets $h:=\{h_{s}\}_{s\in S\setminus\{0\}}\subset\mathbb{C}^{n}$ whose elements are indexed by nonzero elements of $S$.
Proof: We begin by recalling the definition of the groupoid basis $\lfloor s\rfloor$ of the contracted algebra $\mathbb{C}_{0}[S]$ (see Eq. (8)).
@@
where $t\neq 0$. From this expression, it follows that
$\lfloor s^{-1}\rfloor$ $=\sum_{\begin{subarray}{c}t\in S\\ t\leq s\end{subarray}}\mu(t^{-1},s^{-1})t^{-1}=\sum_{\begin{subarray}{c}t\in S\\ t\leq s\end{subarray}}\mu(t^{-1}t,s^{-1}s)t^{-1}=\sum_{\begin{subarray}{c}t\in S\\ t\leq s\end{subarray}}\mu(t,s)t^{-1}.$ (74)
In the last two equalities we have used the fact that $\mu(t,s)=\mu(tt^{-1},ss^{-1})=\mu(t^{-1}t,s^{-1}s)$ (See [13]). Now, with $s,s^{\prime}\neq 0$, we have
@@
Now, by defining a family $h^{0}:=\{h^{0}_{t}\}$ consisting of elements $h^{0}_{t}:=\sum_{\begin{subarray}{c}s\in S\\ t\leq s\end{subarray}}\mu(t,s)h_{s}$ (which can be inverted as $h_{s}=\sum_{\begin{subarray}{c}t\in S\\ t\geq s\end{subarray}}h^{0}_{t}$), we can rewrite the above equation as
@@
for all $\mathcal{D}$-classes $D_{k}$ and all subsets $h:=\{h_{s}\}_{s\in D_{k}}\subset\mathbb{C}^{n}$ whose elements are indexed by the elements of $D_{k}$.
Proof: If $ss^{-1}=s^{\prime}s^{\prime-1}$, then $s$ and $s^{\prime}$ belong to the same $\mathcal{D}$-class. Consequently, if $s$ and $s^{\prime}$ lie in different $\mathcal{D}$-classes, then one has $ss^{-1}\neq s^{\prime}s^{\prime-1}$. The multiplication rule for the groupoid basis (See Eq. (10) implies that $\lfloor s^{-1}\rfloor\lfloor s^{\prime}\rfloor=0$ whenever $ss^{-1}\neq s^{\prime}s^{\prime-1}$. As a result, the matrix $\Phi(\lfloor s^{-1}\rfloor\lfloor s^{\prime}\rfloor)]$ is block diagonal with the blocks indexed by the $\mathcal{D}$-classes of $S$. Therefore, $\Phi$ is positive definite if and only if, for every $\mathcal{D}$-classes $D_{k}$,
@@
for all subsets $h:=\{h_{s}\}_{s\in D_{k}}\subset\mathbb{C}^{n}$ whose elements are indexed by the elements of $D_{k}$. This completes the proof. $\Box$
In the following proposition, we generalize the Schur orthogonality relation to finite inverse semigroups. This result plays a crucial role in the proof of Bochner’s theorem for finite inverse semigroup.
Proposition 5 (Schur Orthogonality relation for inverse semigroup).
Let $S$ be a finite inverse semigroup, and let $\{D_{k}\}_{k=1}^{n}$ denote its $\mathcal{D}$-classes. Fix a $\mathcal{D}$-class $D_{k}$, and let $\overline{\rho}_{k},\overline{\rho^{\prime}}_{k}\in\mathrm{IrrU}(\mathbb{C}_{0}[S])$, where $\mathrm{IrrU}(\mathbb{C}_{0}[S])$ denotes a complete set of inequivalent irreducible representations of $\mathbb{C}_{0}[S]$ induced by a complete set of inequivalent irreducible unitary representations of the maximal subgroups of $S$, as described in Theorem J. Let $a,b,c,d\in D_{k}$ be idempotents. Then
@@
where $\overline{(\cdot)}$ denotes complex conjugation, and $((\overline{\rho}_{k}(\lfloor s\rfloor))_{a,b})_{i,j}$ is the $(i,j)$-th entry of the $(a,b)$-th block of $\overline{\rho}_{k}(\lfloor s\rfloor)$.
Proof: The action of the induced representation $\overline{\rho}_{k}$ on the groupoid basis element $\lfloor s\rfloor$ is given by
@@
Notice that $u:=p^{-1}_{\mathrm{ran}(s)}\,s\,p_{\mathrm{dom}(s)}\in G_{e_{k}}$. Now, from Eq. (81), it follows immediately that
$\overline{((\overline{\rho}_{k}(\lfloor s\rfloor))_{a,b})_{i,j}}=\overline{(\rho_{k}(u))_{i,j}}\delta_{a,\mathrm{ran}(s)}\delta_{b,\mathrm{dom}(s)}.$
Consequently,
$\sum_{s\in D_{k}}((\overline{\rho}_{k}(\lfloor s\rfloor))_{a,b})_{i,j}\overline{((\overline{\rho^{\prime}}_{k}(\lfloor s\rfloor))_{c,d})_{k,l}}=\sum_{s\in D_{k}}(\rho_{k}(u))_{i,j}\delta_{a,\mathrm{ran}(s)}\delta_{b,\mathrm{dom}(s)}\overline{(\rho^{\prime}_{k}(u))_{k,l}}\delta_{c,\mathrm{ran}(s)}\delta_{d,\mathrm{dom}(s)}.$
Next, observe that the map $\vartheta:D_{k}\to M_{r_{k}}(G_{e_{k}})$ defined by
$s\mapsto(p^{-1}_{\mathrm{ran}(s)}\,s\,p_{\mathrm{dom}(s)})E_{\mathrm{ran}(s),\mathrm{dom}(s)}$
is one-to-one correspondence with the inverse given by
$gE_{e,f}\mapsto p_{e}gp_{f}^{-1}.$
Hence any $s\in D_{k}$ can be written as $s=p_{e}gp_{f}^{-1}$ for some pair of idempotents $(e,f)\in D_{k}$ and $g\in G_{e_{k}}$. It should be noted that $\mathrm{ran}(p_{e}gp_{f}^{-1})=e$ and $\mathrm{dom}(p_{e}gp_{f}^{-1})=f$ for all $g\in G_{e_{k}}$. A necessary condition for the terms in the summation of Eq. (82) to be non-vanishing is $a=\mathrm{ran}(s)$ and $b=\mathrm{dom}(s)$. So, the possible non-vanishing contributions come from summing over those $s\in D_{k}$ that have the form $s=p_{a}gp_{b}^{-1}$. With this observation, we have
@@
Now, applying Schur orthogonality relation [42, 43, 44] for the inequivalent irreducible unitary representations $\rho_{k}$ and $\rho^{\prime}_{k}$ of the maximal subgroup $G_{e_{k}}$ in Eq. (83), and using the fact that $d_{\bar{\rho}_{k}}=r_{k}d_{\rho_{k}}$, where $d_{\bar{\rho}_{k}}$ represents the dimension of the induced representation $\overline{\rho}_{k}$, we have
@@
Before proving Bochner’s theorem for finite inverse semigroups, we recall from Eq. (47) that for $s\in D_{k}$, the induced representation is given by
$\overline{\rho}_{k}(\lfloor s\rfloor)=E_{\mathrm{ran}(s),\mathrm{dom}(s)}\otimes\rho_{k}(p^{-1}_{\mathrm{ran}(s)}\,s\,p_{\mathrm{dom}(s)}),$
while $\overline{\rho}_{k}(\lfloor s\rfloor)=0$ if $s\notin D_{k}$. Let $\rho_{k}$ be a unitary representation of the maximal subgroup $G_{e_{k}}$. Then by taking adjoint yields
$(\overline{\rho}_{k}(\lfloor s\rfloor))^{\dagger}=E_{\mathrm{dom}(s),\mathrm{ran}(s)}\otimes\rho_{k}(p^{-1}_{\mathrm{dom}(s)}\,s^{-1}\,p_{\mathrm{ran}(s)}).$
By Eq. (59), this can be written equivalently as
$(\overline{\rho}_{k}(\lfloor s\rfloor))^{\dagger}=\overline{\rho}_{k}(\lfloor s^{-1}\rfloor).$
This will be used in proof Bochner’s theorem for finite inverse semigroup.
Theorem 4 (Bochner’s theorem for finite inverse semigroups).
Let $S$ be a finite inverse semigroup, and let $\Phi:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ be a linear map. Then the linear map $\widetilde{\Phi}:\mathbb{C}_{0}[S]\to M_{n}(\mathbb{C})$ defined by $\widetilde{\Phi}(\lfloor s\rfloor)=\sum_{\begin{subarray}{c}t\in S\\ t\geq s\end{subarray}}\Phi(t)$ is positive definite if and only if the Fourier transform $\widehat{\Phi}(\overline{\rho}_{k})$ is positive semidefinite for all $\overline{\rho}_{k}\in\mathrm{IrrU}(\mathbb{C}_{0}[S])$, where $\mathrm{IrrU}(\mathbb{C}_{0}[S])$ denotes the complete set of inequivalent irreducible representations of $\mathbb{C}_{0}[S]$ induced by a complete set of inequivalent irreducible unitary representations of the maximal subgroups of $S$.
Proof: We begin by recalling the inversion formula for finite inverse semigroups. For $s\in D_{k}$, it is given by
$\widetilde{\Phi}(\lfloor s\rfloor)$ $=\frac{1}{r_{k}|G_{e_{k}}|}\sum_{\sigma\in\mathrm{IrrU}(\mathbb{C}_{0}[S])}d_{\sigma}\operatorname{tr}_{1}\big[(\sigma(\lfloor s^{-1}\rfloor)\otimes\mathbb{I})\widehat{\Phi}(\sigma)\big].$ (85)
Since $\sigma\in\mathrm{IrrU}(\mathbb{C}_{0}[S])$, we have $\sigma(\lfloor s^{-1}\rfloor)=(\sigma(\lfloor s\rfloor))^{\dagger}$. Therefore, the inversion formula can be rewritten as
@@
Let $s,t\in D_{k}$ be such that $s^{-1}s=tt^{-1}$. For such a pair of $s$ and $t$, it follows from Eq. (86) that